# Using Explainable AI for EEG-based Reduced Montage Neonatal Seizure Detection

### v2

@@ VI CONCLUSION
Although neonatal seizure detection is a challenging task even for experienced professionals, by introducing well-defined signal processing techniques and deep learning models, we may attempt to make the task less challenging. By leveraging explainable AI and a seizure detection probability distribution, not only experienced professionals, but less experienced professionals can gain the skillset to diagnose seizure events accurately and provide prompt management. To this end, we have presented an efficient, reliable, and unique deep learning architecture built upon a CNN encoder to extract temporal features, a GAT network to extract spatial features and a binary classification head. On average, it takes only 62 ms to detect seizures in a 12-second EEG epoch on the CPU, with even faster processing (32 ms) on the GPU. Beyond its ability to detect seizures in real-time with a reduced montage, this model offers the unique advantage of real-time interpretability. This allows for quick and insightful analysis. A modified version of Grad-cam is employed to explain the model’s binary class output demonstrating which channels and time windows have been looked at by the model when a seizure is detected. The reduced montage employs only 9 electrodes making it easy to prepare the subject for testing and increasing the patient comfortability.
@@ ACKNOWLEDGEMENT
The authors would like to extend their gratitude to the department for allowing them to access the university computational resources to carry out this research successfully, which were funded by the Accelerating Higher Education Expansion and Development (AHEAD) Operation of the Ministry of Higher Education of Sri Lanka, funded by the World Bank. The authors would also like to thank the staff of The Lady Ridgeway Hospital in Colombo and Department of Paediatrics, Faculty of Medicine, University of Colombo for their dedication and support to Neonatal Care in Sri Lanka.

### v3

@@ VI Conclusion
Although neonatal seizure detection is a challenging task even for experienced professionals, by introducing well-defined signal processing techniques and deep learning models, we may attempt to make the task less challenging. By leveraging explainable AI and a seizure detection probability distribution, not only experienced professionals, but less experienced professionals can gain the skillset to diagnose seizure events accurately and provide prompt management. To this end, we have presented an efficient, reliable, and unique deep learning architecture built upon a CNN encoder to extract temporal features, a GAT network to extract spatial features, and a binary classification head. On average, it takes only $62$ ms to detect seizures in a $12$-second EEG epoch on the CPU, with even faster processing ($32$ ms) on the GPU. Beyond its ability to detect seizures in real time with a reduced montage, this model offers the unique advantage of real-time interpretability. This allows for quick and insightful analysis. A modified version of Grad-cam is employed to explain the model’s binary class output demonstrating which channels and time windows have been looked at by the model when a seizure is detected. The reduced montage employs only $9$ electrodes making it easy to prepare the subject for testing and increasing the patient comfortability.
@@ Acknowledgement
The authors would like to extend gratitude to Dr. Ranga Rodrigo for providing computational resources, which were funded by the Accelerating Higher Education Expansion and Development (AHEAD) grant. The authors would also like to thank the staff of The Lady Ridgeway Hospital in Colombo and the Department of Paediatrics, Faculty of Medicine, University of Colombo for their dedication and support to Neonatal Care in Sri Lanka.